The free cash flow(in millions) shown below are forecast by Parker& Sons.If the weighted average cos of capital(WACC) is 11% and FCF is expected to grow at a rate of 5% after 2years.What is the year 0(Zero) value of operations in millions?
Assume that the ROIC isexpected to remain constant in Year 2 and beyond.(and do not make any half year adjustments)

Year            1              2
FCF            -$50       $100
Answer #1

WACC = 11%

g = growth rate = 5%

FCF in Year 1 = -$50 million

FCF in Year 2 = $100 million

FCF in Year 3 = $100 million * (1+5%) = $105 million

FCF in Year 2 for the years 3 and thereafter = FCF in Year 3 / (WACC - growth rate)

= $105 million / (11% - 5%)

= $1,750 million

Value of Operations in Year 0 = Present value of operations

= [FCF in Year 1 / (1+r)^1] + [FCF in Year 2 / (1+r)^2] + [FCF in Year 2 for years thereafter / (1+r)^2]

= [-$50 million / (1+11%)^1} + [$100 million / (1+11%)^2] + [$1,750 million / (1+11%)^2]

= [-$50 million /1.11] + [$100 million / 1.2321] + [$1,750 million / 1.2321]

= -$45.045045 + $81.162243 + $1,420.33926

= $1,456.45646

Therefore, value of Operations in Year 0 is $1,456.46
Answer #2

FCF1:       -$50

FCF2:      $100

g:               5%

WACC:    11%

First, find the horizon, or terminal, value:

HV2 = FCF2(1 + g)/(WACC – g) = $100(1.05)/(0.11 – 0.05) = $1,750.00

Then find the PV of the free cash flows and the horizon value:

Value of operations = -$50/(1.11) + ($100 + $1,750)/(1.11)2 = $1,456


answered by: Allen